OpenStudy (anonymous):

Dmitri needs 6 liters of a 42% solution of sulfuric acid for a research project in molecular biology. He has two supplies of sulfuric acid solution: one is an unlimited supply of the 54% solution and the other an unlimited supply of the 36% solution. How many liters of each solution should Dmitri use?

OpenStudy (btaylor):

make a system of equations, where A is the 54% and B is the 36%. A + B = 6 <-- describing the total volume .54A + .36 B = (.42)(6) <-- describing the sulfuric acid content now, solve the first for A, and you get A = 6-B So then plug the value for A in, to get .54(6-B)+.36B = 2.52 Can you solve it from here?

OpenStudy (anonymous):

I solved it myself, but in a different way. I got 54% solution: 4 liters; 36% solution: 2 liters?